OpenStudy (anonymous):

Find the x- and y-intercepts of the graph of the equation x2 + y2 = 36

OpenStudy (anonymous):

let x = 0 to find points where y^2 = 36; then let y = 0 to find x^2 = 36 ....

OpenStudy (anonymous):

huh?

myininaya (myininaya):

To find y-intercepts, you set x=0 and solve for y \[0^2+y^2=36 => y^2=36\] You will find 2 y intercepts. Now to find x-intercepts, you set y=0 and solve for x \[x^2+0^2=36=> x^2=36 \] You will find 2 x intercepts. ---------------------------This is what Broken Fixer was saying. ------------------ Now if you realize this is a circle centered at (0,0) with radius 6. You could also find the intercepts just by counting up 6 from (0,0), counting down from (0,0) 6 times, counting to the left of (0,0) 6 times, and counting from the right of (0,0) 6 times.